The battery provides V:52 volts. The capacitors are 1:5 μ F, C2:5pE and C3-4 μ F. Calculate the voltage across capacitor C3. ▼ PAPER SOLUTION Solve the problem on paper first, including all four IDEA steps. Have you finished your paper solution already? oyes no INTERPRET Identify all of the true statements. A. The charge on C2 is equal to the charge on C3 because they are connected in parallel. B. The pair C2 and C3 can be replaced by a single effective capacitor with known capacitance. C. The voltage across C1 is equal to the voltage across C2 because the capacitances are equal D. The charge on C1 is equal to the charge on C2 because the capacitances are equal. E.The voltage across C2 is equal to the voltage across C3 because they are connected in parallel F None of the above.
Answer #1

C2 and C3 are in parallel

C23 =5+4 =9 uF

C1 and C23 are in series ,Therefore equivalent capacitance is

1/C =1/9 +1/5

C=3.2143 uF

Total charge

Q=CV =(3.2143)(52) =167.143 uC

Voltage across capacitor C3 is

V3 = Q/C23 =167.14/9

V3=18.57 Volts

---------------------------------------------------------

  • The pair C2 and C3 can be replaced by a single effective capacitor with known capacitance
  • The Voltage across C2 is equal to the voltage across C3 because they are connected in parallel